# LLM Inference Unveiled: Survey and Roofline Model Insights

### v5

@@ Abstract
The field of efficient Large Language Model (LLM) inference is rapidly evolving, presenting a unique blend of opportunities and challenges. Although the field has expanded and is vibrant, there hasn’t been a concise framework that analyzes the various methods of LLM Inference to provide a clear understanding of this domain. Our survey stands out from traditional literature reviews by not only summarizing the current state of research but also by introducing a framework based on roofline model for systematic analysis of LLM inference techniques. This framework identifies the bottlenecks when deploying LLMs on hardware devices and provides a clear understanding of practical problems, such as why LLMs are memory-bound, how much memory and computation they need, and how to choose the right hardware. We systematically collate the latest advancements in efficient LLM inference, covering crucial areas such as model compression (e.g., Knowledge Distillation and Quantization), algorithm improvements (e.g., Early Exit and Mixture-of-Expert), and both hardware and system-level enhancements. Our survey stands out by analyzing these methods with roofline model, helping us understand their impact on memory access and computation. This distinctive approach not only showcases the current research landscape but also delivers valuable insights for practical implementation, positioning our work as an indispensable resource for researchers new to the field as well as for those seeking to deepen their understanding of efficient LLM deployment. The analyze tool, LLM-Viewer, is open-sourced.
@@ 1 Introduction
Numerous methods have been developed to address the challenges of deploying LLM. The field of efficient LLM inference has grown exponentially in the last two years, presenting both opportunities and challenges. While the burgeoning volume of research demonstrates the field’s vibrancy, it can inadvertently mask key trends and slow advancements. A critical gap in existing literature is the absence of a systematic and practical framework for unified analysis and comprehensive solution development. To bridge this gap, our work offers a comprehensive overview of the current state of research in efficient LLM inference, with a unique focus on its practice-driven characteristics. Diverging from traditional literature reviews, our work not only discusses existing research but also introduces a specifically developed roofline model. This model is designed to analyze bottlenecks in LLM deployments, a crucial step we believe is vital for practical application and optimization as shown in Figure 1. Our work is the first, to our knowledge, that provides such a tool for analyzing the intricacies of inferring LLMs on real devices, systematically collating the latest advancements in efficient LLM inference. We delve deep into deployment challenges, particularly emphasizing inference efficiency. Our discussion spans various areas, including weight optimization, decoding algorithm refinement, system-level and hardware-level enhancements, as illustrated in Figure 2. While there are concurrently related surveys in this domain, such as [Zhu et al., , 2023] on LLM compression and [Miao et al., 2023a, ], [Ding et al., , 2023] and [Wang et al., , 2024] on holistic LLM serving, our work stands out by incorporating a roofline model analysis. This unique aspect allows for a more comprehensive and nuanced understanding of the challenges and solutions in efficient LLM inference, setting our work apart in its practical applicability and thoroughness.
Structure. In this paper, we first discuss the foundations of LLMs, particularly their inference and deployment considerations (Sec.2.1-2). In contrast to traditional survey papers only focusing on literature reviews, we develop a roofline model to practically analyze the bottleneck of deploying LLMs (Sec.2.3-4). In the context of huggingface transformer [^1] https://huggingface.co/docs/transformers/index, the most popular LLM codebase, our roofline model can be used to analyze the deployment of any LLM architecture on any hardware platform, as shown in Figure 1. Importantly, we open-source this roofline models, which can benefit the LLM community. We believe this is an important step to democratize LLM. For the literature review, this survey categorizes strategies for improving LLM inference efficiency into four main areas: Parameter Reduction (Sec.3), Fast Decoding Algorithm Design (Sec.4), Compiler/System-Level Optimization (Sec.5), and Hardware-Level Optimization (Sec.6), providing a comprehensive framework for addressing the complexities of efficient LLM deployment.
@@ 2 Delve into LLM Inference and Deployment
In this chapter, we will delve into the details of how LLMs inference and explore the key considerations in deploying them. By the end of this chapter, you will gain a clear understanding of the operational aspects of LLMs.
@@ 2.1 LLM Inference
2.2 Considerations on Deployment
In the realm of large model inference and deployment, three key factors demand our attention: the computation capability of the hardware, the size of hardware memory, and the bandwidth of that memory. These elements intertwine, playing a pivotal role in the efficiency and success of deploying large models.
Computation Capability: The heart of the matter lies in the computation capability of the hardware. First off, the hardware has to be able to handle the data flow for operations like matrix multiplication and non-linear functions (such as softmax). Also, it needs to support the data types used in these operations, like the common ones - FP16 and FP32, or even INT8 if the model is quantized. If the hardware can handle all the operations needed by the LLMs, then there’s a chance it can be deployed. And don’t forget, the hardware needs enough computational power, meaning it should have plenty of processing units (such as CUDA cores in GPU). We can measure the computational capability by peak computational performance (operations per second, OPS).
Memory Size: For LLMs, having sufficient memory is crucial to hold the weight of the model’s parameters and intermediate results during inference. LLMs’ weights can be several gigabytes (GB) to even over a hundred GB for really big models. The KV cache changes based on how much text you’re working with and the size of the group of tasks you’re doing (batch size). If you’re dealing with a lot of text or have a big batch size, the KV cache can take up several GBs to tens of GBs of space.
Memory Bandwidth: Memory bandwidth refers to the rate at which data moves between the memory and the computation units. When data transfer is slow, the computation units have to wait idly for the data to arrive. This can pose a significant bottleneck for LLMs. In the next section, we will delve into a detailed explanation of this issue.
Choosing the right hardware for LLMs isn’t a one-size-fits-all situation because of the difference of various tasks. The differences in tasks mainly come down to batch size, the ratio between the prefill and decode stages, and the length of the sequences. In the next section, we will take both the hardware and tasks into consideration and introduce how to evaluate the efficiency of inference.
@@ 2.3 Roofline Model
Assessing the efficiency at which LLMs deploy onto specific hardware involves a comprehensive consideration of both hardware and model characteristics. To conduct this evaluation, we employ a technique known as the Roofline model. The Roofline model serves as an effective theoretical framework to assess the potential performance of deploying a model on particular hardware.
@@ 3.1.2 Quantization for Compressing Pre-trained LLMs
To achieve even lower-bit quantization (e.g., below 2-bit), QuIP [Chee et al., , 2023] introduces an innovative approach that accounts for the even distribution of weight magnitudes and the significance of accurately rounding directions unaligned with coordinate axes. QuIP comprises an adaptive rounding procedure that minimizes a quadratic proxy objective, essential for optimizing the quantization process. Additionally, it employs efficient pre- and post-processing techniques that ensure weight and Hessian incoherence through multiplication by random orthogonal matrices, crucial for maintaining quantization effectiveness. Further advancing PTQ methods, Li et al., 2023a are inspired by the observation that aligning the quantized activation distribution with its floating-point counterpart can restore accuracy in LLMs. Their proposed ’Norm Tweaking’ strategy involves a meticulous calibration data generation process and a channel-wise distance constraint. This approach updates the weights of normalization layers, leading to enhanced generalization capabilities. [Shang et al., , 2023] propose partial-binarized LLM (PB-LLM) by introducing binarization [Hubara et al., , 2016] into LLM quantization to push weight quantization under 2 bits. Following PB-LLM, BiLLM [Huang et al., , 2024] pushes weight quantization to almost 1 bit.
@@
Figure 11: Inference time of decode stage for different quantization settings on Llama-2-13b. (Batch size=1).
@@ 3.3.1 White-Box Knowledge Distillation
In white-box distillation, the internal workings and architecture of the teacher model are fully accessible and utilized during the distillation process. This transparency allows the student model to learn not just the output of the teacher model but also its internal representations and decision-making processes. MiniLLM [Gu et al., , 2023] critiques the limitations of standard knowledge distillation objectives and suggests that reverse Kullback-Leibler divergence is more effective for capturing the complexity of generative tasks, which can enhance the student model’s response quality and reliability. MiniLLM also introduces single-step regularization, teacher-mixed sampling, and length normalization to address challenges in training, thus demonstrating great performance potential for distilling LLMs on the standard benchmarks. In contrast to MiniLLM, GKD [Agarwal et al., , 2023] presents a more straightforward and stable method. It aligns more with supervised training by avoiding backpropagation through the student model’s sampling. Instead of using predetermined output sequences, GKD trains the student model on its own created sequences, utilizing the teacher’s probabilities as guidance, which leads to notable improvements in the student’s performance. Homotopic distillation [Liang et al., 2023a, ] aims to facilitate the alignment of the student model’s predictions with those of the teacher model across extensive open-domain data. It involves starting the student model with the teacher model’s configuration and progressively reducing the student model’s neurons to reach a specified model complexity. Furthermore,  Liang et al., 2023b present a layerwise distillation approach that involves creating unique task-aware filters for each layer of teacher and student models. These filters, essentially neural networks equipped with task-specific heads, are designed to distill and capture the predictive knowledge from the hidden layers of the respective models. AD-KD [Wu et al., 2023b, ] analyzes the teacher model’s token-level rationale using integrated gradients and transfers attributional knowledge to the student model, which enables the student model to imitate the teacher’s underlying reasoning, not just its behaviors.
@@ 3.3.2 Black-Box Knowledge Distillation
Contrary to white-box distillation, black-box distillation does not require access to the internal mechanisms of the teacher model. Instead, it focuses on replicating the output behavior of the teacher model. The student model learns solely from the input-output pairings produced by the teacher, without any insight into its internal operations. Multitask-ICT [Huang et al., , 2022] introduces in-context learning distillation, which merges the objectives of in-context learning with those of language modeling, intending to distill into smaller models both the capability to comprehend in-context examples and the knowledge required for specific tasks. LaMini-LM [Wu et al., 2023a, ] creates a set of 2.58 million instructions and employs GPT-3.5 Turbo to produce responses to these instructions. Subsequently, it uses these instructions as a basis to fine-tune a range of student models. Similarly proceeding from creating examples,  Sahu et al., [2023] proposes PromptMix, which involves a two-step method based on prompting to create labeled examples for text classification. In PromptMix, the borderline examples can enhance the knowledge transfer from teacher models like GPT-3.5 to student models. In contrast to the traditional unidirectional knowledge distillation, Lion [Jiang et al., , 2023] introduces an adversarial distillation framework, which encourages the teacher model to identify ”hard” instructions and subsequently generate new ”hard” instructions for the student model, resulting in a dynamic three-step adversarial cycle.
@@ 4 Algorithmic Methods for Fast Decoding
Large Language Models (LLM) have achieved astonishing performance in various text generation tasks such as Machine Translation, Article Summarization, and Question Answering. To complete these tasks, popular LLMs typically contain the decoder structure that generates tokens one after another, following an autoregressive relation with all preceding tokens. During the decoding of every token, decoder weights have to be repeatedly loaded into memory. Because of the autoregressive nature, the memory loading is innately sequential and cannot be easily parallelized. As the parameter size of LLM becomes colossal, the decoding process becomes heavily memory-bound (de Jong et al., [2023]) and experiences low hardware utilization, leading to severely long latency (Kim et al., 2023d ). Text generation tasks are the fundamental building block for extremely popular real-world applications such as ChatBot (OpenAI et al., [2023]) which demands swift even real-time responses. Finding ways to reduce LLM inference latency is now an urgent problem.
This section focuses on the discussion of prior efforts to reduce the LLM inference latency from an algorithm perspective. We are aware of the ample previous attempts from the system and hardware perspectives and defer the discussion of them in Sections 5 and 6. Also, model compression techniques introduced in Section 3 can directly reduce the parameter size of the LLM and decrease the inference latency, as LLM inference is memory-bound. However, these techniques inevitably lead to performance degradation to tradeoff the model size shrinkage, which is potentially unacceptable for certain real-world applications. This section looks at prior works that reduce LLM inference latency but need to store the entire LLM in the machine storage to preserve the text generation quality. Specifically, this section intends to develop the discussion from two directions:
@@
  • In section 13, for every single forward propagation of the LLM (fixed #parameters used) how to decode the maximum number of tokens.
@@ 4.1 Minimum Parameter Used Per Token Decoded
Interestingly, Simoulin and Crabbé, [2021] has shown that although language models tend to have a huge number of parameters, not all parameters are needed to decode the accurate tokens. LLM inference latency can be reduced by selecting only a subset of necessary parameters to use (load) per input token and still preserving the decoded token’s accuracy. In this section, we look at input-dependent dynamic weights dropping scheme for LLM from three different perspectives: 4.1.1 looks at early exiting, or dynamically choosing weights in the layer, depth, dimension; 4.1.2 introduces novel methods that dynamically detects sparsity in the width dimension of the LLM, pruning out heads and MLP columns; 4.1.3 present Mixture-of-Experts (MoE) which pretrains a sparse model and chooses the correct experts for the different input during runtime.
@@ 4.1.1 Early Exiting
Early exiting has been a well-explored idea in various network architectures, particularly for the encoder-only models. (Baier-Reinio and Sterck, [2020], Hou et al., [2020], Li et al., [2021], Liu et al., [2020], Liu et al., [2022], Schwartz et al., [2020], Stickland and Murray, [2019], Xin et al., [2020], Zhou et al., [2020], Zhu, [2021], Schuster et al., [2021]). However, the LLMs that specialize in continuous token generation at the sequence level rely particularly on the decoder architecture. Early exiting for decoder architecture requires consistency and quality retaining on the sequence level where each token depends on the previous tokens, which are considerations lacking in the previous abundant encoder-only early-exiting literature. The decoder contains layers of identical structure. Benefiting from this trait, the output hidden states of every layer can be used to pass in the LM Head to get a probability distribution prediction of the next token decoded. Geva et al., [2022] and Simoulin and Crabbé, [2021] observe that for some tokens, the hidden states saturate during the intermediate layers. In other words, early exiting in the middle would, for some tokens, output the correct top-1 prediction as running through the full model. This observation lays the basis for the success of decoder early exiting methods.
@@
Modeling the Confidence of Saturation CALM Schuster et al., [2022] studies three different ways to output the confidence score to exit: the softmax response, or the difference between the top two values after the softmax; the saturation of hidden states, or the cosine similarity between the current layer’s hidden states with the last layer; the output of a linear classifier inserted to every layer. The linear classifier is trained by simply using a cross-entropy loss to align MLP output when inputting the hidden states with whether the top-1 token decoded as exiting on the current layer matches the top-1 token decoded of the full model. The experiments presented suggest that despite not being the most accurate predictor, the classifier method reaches the optimal trade-off between additional FLOPs overhead with prediction accuracy on score generating. Building up from CALM, Bae et al., [2023] observed that when consistently exiting from shallow layers will result in an abnormally long length. Also, the confidence score computation on every layer injects high overhead and diminishes the benefit of early exiting. Therefore, it proposes to only have two choices for early exiting: either exit from the so-called ”shallow module” or a group of shallow layers, or go all the way to the full model, or ”deep module”, drastically reducing the number of classifiers needed inside the model. Such design enables it to achieve more speedup than CALM, reaching 2x for certain tasks. On the other hand, ConsistentEE Zeng et al., [2023] proposes a different method to predict when to exit. It uses an RL policy network that is iteratively trained with the per-layer output classifier head. The policy networks are trained with the goal of balancing the optimization of both efficiency (the early layer receives rewards) and accuracy (the reward function has a term that is the early exit output CE loss).
@@
Early Exit Criteria CALM Schuster et al., [2022] proposes a distribution-free calibration technique that uses the fixed sequence testing procedure (Family-wise Error Rate procedure) to output the suitable threshold. The threshold is exponentially decreasing to allow more aggressive exiting for tokens later in the sequence. Bae et al., [2023], on the other hand, observes that the pattern of confidence criteria resembles a beta distribution and uses the on-the-fly data to update a beta distribution model through MLE and use such probability model to guide its decision. Zeng et al., [2023] bypasses this issue by letting the policy network directly output the exit decision.
@@
Hidden States Propagation Hidden states of the skipped layers can pose a technical challenge. As shown in the 12 (b), the token position at ”school” exits later than previous tokens. However, the last self-attention layer doesn’t have the previous key-value pairs of the previous early exited tokens. Elbayad et al., [2020] and Schuster et al., [2022] proposes the ”hidden states propagation” technique. For example, the hidden states of token ”Max” at the exited layer $l_{1}$ are stored. When the later token ”school” reaches deeper layer $l_{2}$, the hidden state for ”Max” is copied for all layers between $l_{1}$ and $l_{2}$, and the key-value pairs are then computed on the copied hidden states. Basically, to approximate the deep layer’s hidden state with the ones from the early layer. Later works Bae et al., [2023] and Ding et al., [2023] found that state propagation leads to performance degradation. Since LLM inferences are dominated mostly by memory loading, computation is relatively ”free”. These two methods proposed to recompute the later hidden states directly on the fly. Chen et al., 2023b proposes to run the full large model in parallel to the early exit stream to efficiently parallel the computation of the missing kv cache. Din et al., [2023] conducts a systematic study on using a linear network to jump across layers for transformer architecture and shows that linear layers can be added to effectively bridge the performance gap between directly copying and computing the hidden states with low memory and compute cost. SkipDecode Corro et al., [2023] chooses an aggressive approach to prioritize the speedup and relax the performance preservation goal. By utilizing the observation that a token coming later in the same sequence on average requires less number of layers to decode the correct tokens, it completely bypasses the need for state propagation by forcing the maximum layer used to be monotonically decreasing for deeper positions. Besides, SkipDecode also introduces fixed exit points to optimize for batched early exit.
@@
Output Classifier Training When exiting from intermediate layers, the intermediate hidden states need to go through an output classifier head to output prediction of the next token probability distribution. The output classifier can either be shared as shown in Figure 12 or per-layer independent. These classifiers are usually trained to better adapt to the early exiting pattern. Elbayad et al., [2020] proposed to have an average CE loss of all layers to be the training loss of the classifier. On the other hand, Schuster et al., [2022] uses a weighted average where weights increase as the layer number increases, assigning more contribution to deeper layers. Bae et al., [2023] introduces a dynamic knowledge distillation loss which dynamically assigns the ”shallow module” a suitable hidden state from the ”deep module”. Both Rotem et al., [2023] and Ji et al., [2023] find a ”conflicting gradient” issue when joint training with the same loss across all models: Rotem et al., [2023] detects the gradient conflict between early and later layers of language models, while Ji et al., [2023] spots the ”orthogonal gradient” between the objective to improve semantic awareness and the objective to improve early exciting decision. Both methods propose adding an additional block of parameters and iterative training to alleviate the issue.
Besides the above-mentioned perspectives, Chen et al., 2023b studies system-level optimization techniques to efficiently run LLM early exit under the 3D parallelism setting.

### v6

@@ Abstract
The field of efficient Large Language Model (LLM) inference is rapidly evolving, presenting a unique blend of opportunities and challenges. Although the field has expanded and is vibrant, there hasn’t been a concise framework that analyzes the various methods of LLM Inference to provide a clear understanding of this domain. Our survey stands out from traditional literature reviews by not only summarizing the current state of research but also by introducing a framework based on Roofline model for systematic analysis of LLM inference techniques. This framework identifies the bottlenecks when deploying LLMs on hardware devices and provides a clear understanding of practical problems, such as why LLMs are memory-bound, how much memory and computation they need, and how to choose the right hardware. We systematically collate the latest advancements in efficient LLM inference, covering crucial areas such as model compression (e.g., quantization), algorithm improvements (e.g., speculative decoding), and both system and hardware-level enhancements (e.g., operator fusion). Our survey stands out by analyzing these methods with Roofline model, helping us understand their impact on memory access and computation. This distinctive approach not only showcases the current research landscape but also delivers valuable insights for practical implementation, positioning our work as an indispensable resource for researchers new to the field as well as for those seeking to deepen their understanding of efficient LLM deployment. The analyze tool, LLM-Viewer, is open-sourced.
@@ 1 Introduction
Numerous methods have been developed to address the challenges of deploying LLM. The field of efficient LLM inference has grown exponentially in the last two years, presenting both opportunities and challenges. While the burgeoning volume of research demonstrates the field’s vibrancy, it can inadvertently mask key trends and slow advancements. A critical gap in existing literature is the absence of a systematic and practical framework for unified analysis and comprehensive solution development. To bridge this gap, our work offers a comprehensive overview of the current state of research in efficient LLM inference, with a unique focus on its practice-driven characteristics. Diverging from traditional literature reviews, our work not only discusses existing research but also introduces a specifically developed Roofline model. This model is designed to analyze bottlenecks in LLM deployments, a crucial step we believe is vital for practical application and optimization as shown in Figure 1. Our work is the first, to our knowledge, that provides such a tool for analyzing the intricacies of inferring LLMs on hardware devices, systematically collating the latest advancements in efficient LLM inference. We delve deep into deployment challenges, particularly emphasizing inference efficiency. Our discussion spans various areas, including model compression, decoding algorithm refinement, system-level and hardware-level enhancements, as illustrated in Figure 2. While there are concurrently related surveys in this domain, such as [Zhu et al., , 2023] on LLM compression and [Miao et al., 2023a, ], [Ding et al., , 2023] and [Wang et al., 2024a, ] on holistic LLM serving, our work stands out by incorporating a Roofline model analysis.
In this paper, we first discuss the foundations of LLMs and develop a tool named LLM-Viewer, which uses Roofline model to analyze the bottleneck of deploying LLMs (Sec.2). LLM-Viewer can be used to analyze the deployment of any LLM architecture on various hardware platform, as shown in Figure 1. For the literature review, this survey categorizes strategies for improving LLM inference efficiency into four main areas: Model Compression (Sec.3), Algorithmic Methods for Fast Decoding (Sec.4), Compiler/System-Level Optimization (Sec.5), and Hardware-Level Optimization (Sec.6).
@@ 2.2 Roofline Model
Assessing the efficiency at which LLMs deploy onto specific hardware involves a comprehensive consideration of both hardware and model characteristics. To conduct this evaluation, we employ the Roofline model. The Roofline model serves as an effective theoretical framework to assess the potential performance of deploying a model on particular hardware.
@@ 3.1.2 Quantization for Compressing Pre-trained LLMs
To achieve even lower-bit quantization (e.g., below 2-bit), QuIP [Chee et al., , 2023] introduces an innovative approach that accounts for the even distribution of weight magnitudes and the significance of accurately rounding directions unaligned with coordinate axes. QuIP comprises an adaptive rounding procedure that minimizes a quadratic proxy objective, essential for optimizing the quantization process. Additionally, it employs efficient pre- and post-processing techniques that ensure weight and Hessian incoherence through multiplication by random orthogonal matrices, crucial for maintaining quantization effectiveness. Further advancing PTQ methods, Li et al., 2023a are inspired by the observation that aligning the quantized activation distribution with its floating-point counterpart can restore accuracy in LLMs. Their proposed ’Norm Tweaking’ strategy involves a meticulous calibration data generation process and a channel-wise distance constraint. This approach updates the weights of normalization layers, leading to enhanced generalization capabilities. [Shang et al., , 2024] propose partial-binarized LLM (PB-LLM) by introducing binarization [Hubara et al., , 2016] into LLM quantization to push weight quantization under 2 bits. Following PB-LLM, BiLLM [Huang et al., , 2024] pushes weight quantization to almost 1 bit.
@@
Figure 11: Memory Consumption of decode stage for different quantization settings on Llama-2-13b. (Batch size=1).
@@ 3.3.1 White-Box Knowledge Distillation
In white-box distillation, the architecture and weights of the teacher model are fully accessible. This transparency allows the student model to learn not just the output of the teacher model but also its internal representations and decision-making processes. MiniLLM [Gu et al., , 2023] critiques the limitations of standard knowledge distillation objectives and suggests that reverse Kullback-Leibler divergence is more effective for capturing the complexity of generative tasks, which can enhance the student model’s response quality and reliability. MiniLLM also introduces single-step regularization, teacher-mixed sampling, and length normalization to address challenges in training, thus demonstrating great performance potential for distilling LLMs on the standard benchmarks. In contrast to MiniLLM, GKD [Agarwal et al., , 2023] presents a more straightforward and stable method. It aligns more with supervised training by avoiding backpropagation through the student model’s sampling. Instead of using predetermined output sequences, GKD trains the student model on its own created sequences, utilizing the teacher’s probabilities as guidance, which leads to notable improvements in the student’s performance. Homotopic distillation [Liang et al., 2023a, ] aims to facilitate the alignment of the student model’s predictions with those of the teacher model across extensive open-domain data. It involves starting the student model with the teacher model’s configuration and progressively reducing the student model’s neurons to reach a specified model complexity. Furthermore,  Liang et al., 2023b present a layerwise distillation approach that involves creating unique task-aware filters for each layer of teacher and student models. These filters, essentially neural networks equipped with task-specific heads, are designed to distill and capture the predictive knowledge from the hidden layers of the respective models. AD-KD [Wu et al., 2023b, ] analyzes the teacher model’s token-level rationale using integrated gradients and transfers attributional knowledge to the student model, which enables the student model to imitate the teacher’s underlying reasoning, not just its behaviors.
@@ 3.3.2 Black-Box Knowledge Distillation
Contrary to white-box distillation, black-box distillation does not require access to the internal information of the teacher model. Instead, it focuses on replicating the output behavior of the teacher model. The student model learns solely from the input-output pairings produced by the teacher, without any insight into its internal operations. Multitask-ICT [Huang et al., , 2022] introduces in-context learning distillation, which merges the objectives of in-context learning with those of language modeling, intending to distill into smaller models both the capability to comprehend in-context examples and the knowledge required for specific tasks. LaMini-LM [Wu et al., 2023a, ] creates a set of 2.58 million instructions and employs GPT-3.5 Turbo to produce responses to these instructions. Subsequently, it uses these instructions as a basis to fine-tune a range of student models. Similarly proceeding from creating examples,  Sahu et al., [2023] proposes PromptMix, which involves a two-step method based on prompting to create labeled examples for text classification. In PromptMix, the borderline examples can enhance the knowledge transfer from teacher models like GPT-3.5 to student models. In contrast to the traditional unidirectional knowledge distillation, Lion [Jiang et al., , 2023] introduces an adversarial distillation framework, which encourages the teacher model to identify ”hard” instructions and subsequently generate new ”hard” instructions for the student model, resulting in a dynamic three-step adversarial cycle.
@@ 4 Algorithmic Methods for Fast Decoding
LLMs have achieved astonishing performance in various text generation tasks. They typically contain the decoder stage that generates tokens one after another, following an autoregressive relation with all preceding tokens. During the decoding of every token, decoder weights have to be repeatedly loaded into memory. As the parameter size of LLM becomes colossal, the decoding process becomes heavily memory-bound [de Jong et al., , 2023] and experiences low hardware utilization, leading to severely long latency [Kim et al., 2023d, ]. This is particularly problematic in real-world applications like ChatBot, where quick and even real-time responses are crucial. Therefore, there is a strong need to optimize the decoding process to improve performance in such applications.
This section focuses on the discussion of prior efforts to reduce the LLM inference cost from an algorithm perspective. Specifically, this section intends to develop the discussion from two directions:
@@
  • In section 4.2, for every single forward propagation of the LLM (fixed #parameters used) how to decode the maximum number of tokens.
@@ 4.1 Minimum Parameter Used Per Token Decoded
Interestingly, Simoulin and Crabbé, [2021] has shown that although language models tend to have a huge number of parameters, not all parameters are needed to generate the accurate tokens. LLM inference latency can be reduced by selecting only a subset of necessary parameters to use (load) per input token and still preserving the decoded token’s accuracy. In this section, we look at input-dependent dynamic weights dropping scheme for LLM from three different perspectives: 4.1.1 looks at early exiting, or dynamically choosing weights in the layer, depth, dimension; 4.1.2 introduces methods that dynamically detects sparsity in the width dimension of the LLM, pruning out heads and MLP columns; 4.1.3 present Mixture-of-Experts (MoE) which pretrains a sparse model and chooses the correct experts for the different input during runtime.
@@ 4.1.1 Early Exiting
Early exiting (or layer skipping) has been a well-explored idea in various network architectures, particularly for the encoder-only models [Baier-Reinio and Sterck, , 2020, Hou et al., , 2020, Li et al., , 2021, Liu et al., , 2020, 2022, Schwartz et al., , 2020, Stickland and Murray, , 2019, Xin et al., , 2020, Zhou et al., , 2020, Zhu, , 2021, Schuster et al., , 2021]. Early exiting for decoder architecture requires consistency and quality retaining on the sequence level where each token depends on the previous tokens, which are considerations lacking in the previous abundant encoder-only early-exiting literature. The decoder contains layers of identical structure. Benefiting from this trait, the output hidden states of every layer can be used to pass in the LM Head to get a probability distribution prediction of the next token decoded. Geva et al., [2022] and Simoulin and Crabbé, [2021] observe that for some tokens, the hidden states saturate during the intermediate layers. In other words, early exiting in the middle would, for some tokens, output the correct top-1 prediction as running through the full model. This observation lays the basis for the success of decoder early exiting methods.
@@
Modeling the Confidence of Saturation. CALM [Schuster et al., , 2022] studies three different ways to output the confidence score to exit: the softmax response, or the difference between the top two values after the softmax; the saturation of hidden states, or the cosine similarity between the current layer’s hidden states with the last layer; the output of a linear classifier inserted to every layer. The linear classifier is trained by simply using a cross-entropy loss to align MLP output when inputting the hidden states with whether the top-1 token decoded as exiting on the current layer matches the top-1 token decoded of the full model. The experiments presented suggest that despite not being the most accurate predictor, the classifier method reaches the optimal trade-off between additional FLOPs overhead with prediction accuracy on score generating. Building up from CALM, [Bae et al., , 2023] observed that when consistently exiting from shallow layers will result in an abnormally long length. Also, the confidence score computation on every layer injects high overhead and diminishes the benefit of early exiting. Therefore, it proposes to only have two choices for early exiting: either exit from the so-called ”shallow module” or a group of shallow layers, or go all the way to the full model, or ”deep module”, drastically reducing the number of classifiers needed inside the model. Such design enables it to achieve more speedup than CALM, reaching 2x for certain tasks. On the other hand, ConsistentEE Zeng et al., [2023] proposes a different method to predict when to exit. It uses an RL policy network that is iteratively trained with the per-layer output classifier head. The policy networks are trained with the goal of balancing the optimization of both efficiency (the early layer receives rewards) and accuracy (the reward function has a term that is the early exit output CE loss).
@@
Early Exit Criteria. CALM Schuster et al., [2022] proposes a distribution-free calibration technique that uses the fixed sequence testing procedure (Family-wise Error Rate procedure) to output the suitable threshold. The threshold is exponentially decreasing to allow more aggressive exiting for tokens later in the sequence. Bae et al., [2023], on the other hand, observes that the pattern of confidence criteria resembles a beta distribution and uses the on-the-fly data to update a beta distribution model through MLE and use such probability model to guide its decision. Zeng et al., [2023] bypasses this issue by letting the policy network directly output the exit decision.
@@
Hidden States Propagation. Hidden states of the skipped layers can pose a technical challenge. As shown in the 12 (b), the token position at ”school” exits later than previous tokens. However, the last self-attention layer doesn’t have the previous key-value pairs of the previous early exited tokens. Elbayad et al., [2020] and Schuster et al., [2022] proposes the ”hidden states propagation” technique. For example, the hidden states of token ”Max” at the exited layer $l_{1}$ are stored. When the later token ”school” reaches deeper layer $l_{2}$, the hidden state for ”Max” is copied for all layers between $l_{1}$ and $l_{2}$, and the key-value pairs are then computed on the copied hidden states. Basically, to approximate the deep layer’s hidden state with the ones from the early layer. Later works Bae et al., [2023] and Ding et al., [2023] found that state propagation leads to performance degradation. Since LLM inferences are dominated mostly by memory loading, computation is relatively ”free”. These two methods proposed to recompute the later hidden states directly on the fly. Chen et al., 2023b proposes to run the full large model in parallel to the early exit stream to efficiently parallel the computation of the missing kv cache. Din et al., [2023] conducts a systematic study on using a linear network to jump across layers for transformer architecture and shows that linear layers can be added to effectively bridge the performance gap between directly copying and computing the hidden states with low memory and compute cost. SkipDecode Corro et al., [2023] chooses an aggressive approach to prioritize the speedup and relax the performance preservation goal. By utilizing the observation that a token coming later in the same sequence on average requires less number of layers to decode the correct tokens, it completely bypasses the need for state propagation by forcing the maximum layer used to be monotonically decreasing for deeper positions. Besides, SkipDecode also introduces fixed exit points to optimize for batched early exit.
@@
Output Classifier Training. When exiting from intermediate layers, the intermediate hidden states need to go through an output classifier head to output prediction of the next token probability distribution. The output classifier can either be shared as shown in Figure 12 or per-layer independent. These classifiers are usually trained to better adapt to the early exiting pattern. Elbayad et al., [2020] proposed to have an average CE loss of all layers to be the training loss of the classifier. On the other hand, Schuster et al., [2022] uses a weighted average where weights increase as the layer number increases, assigning more contribution to deeper layers. Bae et al., [2023] introduces a dynamic knowledge distillation loss which dynamically assigns the ”shallow module” a suitable hidden state from the ”deep module”. Both Rotem et al., [2023] and Ji et al., [2023] find a ”conflicting gradient” issue when joint training with the same loss across all models: Rotem et al., [2023] detects the gradient conflict between early and later layers of language models, while Ji et al., [2023] spots the ”orthogonal gradient” between the objective to improve semantic awareness and the objective to improve early exciting decision. Both methods propose adding an additional block of parameters and iterative training to alleviate the issue. Besides the above-mentioned perspectives, Chen et al., 2023b studies system-level optimization techniques to efficiently run LLM early exit under the 3D parallelism setting.